OpenStudy (loser66):

Explain me, please. I don't get why they said so. x1 = 0.85779 . Is it not that x1 = 85779/100000 ? the same with others x2,x3,x4... why do they say the polynomial doesn't have rational root?

OpenStudy (loser66):

@xapproachesinfinity help me, kid

Miracrown (miracrown):

Maybe there is not just 0.85779 that could be just the first 5 decimals like say pi is 3.14159...

Miracrown (miracrown):

It must be, or it would be rational. That is the root of what polynomial?

OpenStudy (loser66):

:) It is another topic. The original one is "find the minimal polynomial for sin(2pi/5) I end up with 16t^4-20t^2+5 The last step is proving it is irreducible.

Miracrown (miracrown):

85779/100000 is the same as 0.85779

OpenStudy (loser66):

irreducible in Q. That is it doesn't have a rational root.

OpenStudy (loser66):

I tested by using that tool. It gives me that answer but I don't understand :)

OpenStudy (perl):

you can use rational roots theorem to show that 16t^4-20t^2+5=0 has no rational roots (by testing them)

OpenStudy (zarkon):

just solve it...it is a quadratic in \(t^2\)

Miracrown (miracrown):

What the website is saying is that there rational roots that can be found using the rational root test. It's not saying that 0.58779 is not a rational number

OpenStudy (loser66):

@Miracrown it says: This polynomial has no rational roots that can be found using Rational Root Test.

Miracrown (miracrown):

Yes, so

OpenStudy (loser66):

@Zarkon you mean I let t^2 = x then solve that quadratic to see that there is no rational root?

OpenStudy (zarkon):

yes

OpenStudy (perl):

check the discriminant

OpenStudy (loser66):

I got it. Thank you to all.

OpenStudy (loser66):

Proposition: If \(K\subseteq L\) is a finite field extension, then a) L is algebraic over K b) there are finitely many (algebraic) elements a1,a2,......,a_n in L such that L =K(a1,a2,...,a_n) @Miracrown question follows

OpenStudy (loser66):

If the conclusion of part b is false, show that there exists an infinite sequence a1,a2,....... of elements of L such that \(K\subset K(a_1)\subset K(a_1,a_2)\subset K(a_1,a_2,a_3)\subset........\) Show that this contradicts the finiteness of \(dim_K(L)\)

OpenStudy (loser66):

My question: How many parts do I have for this problem? 1 or 2? the first "show" is to show there does not exist the sequence above, and the second "show" is separate to it, right?